# Semiregular abstract polyhedra with trivial facet stabilizer

### v1

@@ Abstract
Abstract polytopes generalize the face lattice of convex polytopes. An (abstract) polytope is semiregular if its facets are regular and its automorphism group acts transitively on its vertices. In this paper we construct semiregular, facet-transitive polyhedra with trivial facet stabilizer, showing that semiregular abstract polyhedra can have an unbounded number of flag orbits, while having as little as one facet orbit. We interpret this construction in terms of operations applied to high rank regular and chiral polytopes, and we see how this same operations help us construct alternating semiregular polyhedra. Finally, we give an idea to generalize this construction giving examples in higher ranks.
@@ 1 Introduction
Abstract polytopes are combinatorial objects that generalize the face lattice of convex polytopes and other geometric objects. Just as with geometric polytopes, the main approach to study abstract polytopes has been through their symmetry. A flag is typically defined as a set having one face of each rank, and satisfying that all those faces are incident to each other. One can prove that the automorphism group of an abstract polytope acts semiregularly on its flags, and therefore, one can measure how symmetric a polytope is by counting its flag orbits: the less flag orbits the more symmetrical. Regular polytopes are those with exactly one flag orbit, and are by far the most studied class of abstract polytope, having a whole book [10] dedicated to them.
@@
When studying other classes of polytopes, some have studied polytopes with a small number of flag orbits (see for example [6], [12], [16], [17]), but some classes where the number of flag orbits is not restricted have been studied too. Such is the case of semiregular polytopes: those with regular facets and where the automorphism group acts transitively on the vertices. Despite the number of flag orbits being unbounded, most of the research on semiregular polytopes has been focused on constructions where the result has only one or two flag orbits (see [14], for example). This is due to the fact that the examples constructed are hereditary: every symmetry of a facet induces a symmetry of the whole polytope.
@@
In this paper we construct semiregular polyhedra that are face-transitive, that are the opposite of hereditary: No symmetry of a facet induces a symmetry of the whole polyhedron, in other words, our examples have trivial facet stabilizer. This implies that despite having one orbit on vertices and one orbit on faces, the number of orbits on flags in unbounded in our construction.
@@
For our construction is convinient to think of polytopes as a special class of maniplexes: graphs with a proper coloring of their edges satisfying certain conditions. Thinking of polytopes in terms of graphs allows us to use voltage graphs to construct them, and thanks to the results of [7] and [1]. we know conditions on the voltage assignments that tell us when the graphs we obtain represent a polytope.
@@
We will also see that our examples can be interpreted as the result of applying some particular operations to regular and chiral polytopes of high rank. This same operations can be used to obtain alternating semi-regular polyhedra with trivial facet stabilizer (that is, semiregular polyhedra with two face-orbits and where each edge lies between faces in different orbits).
@@
In Section 2 we introduce the basic notions necessary for this paper. Section 3 presents the concept of symmetry type graph, which is used a lot later in the paper. Section 4 presents the non-standard definition of graph used in this paper, together with the concept of voltage graphs and how to construct the derived graph of a voltage graph. In Section 5 we characterize the regular symmetry type graphs that semiregular facet-transitive polyhedra with trivial facet stabilizer may have. Section 6 and Section 7 deal with how to construct polyhedra with the symmetry types determined in Section 5. In Section 8 we interpret our construction in terms of operations. In Section 9 we see how to use these operations to get examples of alternating semiregular polyhedra. Finally, in Section 10 we see how we could generalize some of our constructiions to get semiregular, facet-transitive maniplexes with trivial facet stabilizer in higher rank, although we are not sure yet how to make sure that they are polytopes.
@@ 2 Basic concepts
Abstract polytopes, introduced by Schulte and Danzer in [3], are typically defined as posets with certain properties that make them similar to the face lattice of a convex polytope. However, thanks to [4, Theorem 5.3], we know that they can be equivalently defined as a particular kind of maniplex. For simplicity, we will use this other definition.
An $n$-maniplex (or maniplex of rank $n$) is a connected simple graph $\mathcal{M}$ with a coloring of its edges with colors in $[0,n-1]:=\{0,1,\ldots,n-1\}$ such that
@@
  • given two colors $i$ and $j$ such that $|i-j|>1$, the alternating paths of length 4 using edges of these two colors are closed.
The vertices of a maniplex are referred to as flags. For the purposes of this paper all maniplexes are finite, that is, they have a finite number of flags. Maniplexes are a natural generalization of maps to higher dimmensions. To see other definitions of maniplex and also learn the topological interpretation of this definition see [20].
Given an $n$-maniplex $\mathcal{M}$ and a color $i\in[0,n-1]$, we denote by $\mathcal{M}_{\overline{i}}$ the subgraph of $\mathcal{M}$ obtained by removing the edges of color $i$. The connected components of $\mathcal{M}_{\overline{i}}$ are the $i$-faces of $\mathcal{M}$. The $(n-1)$-faces of an $n$-maniplex are its facets, while the 0-faces and 1-faces are its vertices and edges respectively.
We say that an $i$-face $F$ and a $j$-face $G$ of a maniplex are incident if they have non-empty intersection. If in addition $i\leq j$, we write $F\leq G$. Note that in this way the faces of a maniplex form a poset.
In Figure 1 we see how a hexagon can be interpreted as a 2-maniplex. Note that the vertices of the hexagon correspond to the edges of color 1 of the graph, and the edges of the hexagon correspond to the edges of color 0. Analogously, in Figure 2 we see a triangular prism interpreted as a 3-maniplex. Here the vertices of the prism correspond to cylces with edges of colors 1 and 2 in the graph, the edges of the prism correspond to cycles with edges of colors 0 and 2 in the graph, and the 2-faces of the prism correspond to cycles with edges of colors 0 and 1 in the graph.
@@
Given a maniplex $\mathcal{M}$ we define its dual $\mathcal{M}^{*}$ as the maniplex we get by replacing the edges of color $i$ with edges of color $n-1-i$ for each $i\in[0,n-1]$. Note that the vertices of $\mathcal{M}^{*}$ correspond to the facets of $\mathcal{M}$ and vice-versa. If $\mathcal{M}^{*}$ is isomorphic to $\mathcal{M}$ we say that $\mathcal{M}$ is self-dual.
@@
Path intersection property: Let $\Phi$ and $\Psi$ be two flags of $\mathcal{P}$. If there is a path from $\Phi$ to $\Psi$ that uses only colors in $[0,m]$ and another path from $\Phi$ to $\Psi$ that uses only colors in $[k,n-1]$, then there is a path from $\Phi$ to $\Psi$ that uses only colors in $[k,m]$.
In [4, Theorem 5.3] it is proved that this definition is equivalent to the poset definition given by Shculte and Danzer in  [3]. From here on, whenever we talk about polytopes in this paper we will be thinking about abstract polytopes.
@@
An isomorphism of $n$-maniplexes is just a graph isomorphism that preserves the colors of the edges. Naturally, an automorphism of a maniplex is just an isomorphism onto itself. The automorphism group of a maniplex $\mathcal{M}$ is denoted by $\Gamma(\mathcal{M})$. It is known that the automorphism group of a maniplex acts freely on its flags (the proof is identical to that of [10, Proposition 2A4], which is the particular case of abstract polytopes). We say that a maniplex $\mathcal{M}$ is regular (some authors prefer reflexive) if $\Gamma(\mathcal{M})$ acts transitively (and therefore regularly) on the flags of $\mathcal{M}$.
A maniplex is called semiregular if it is vertex-transitive (that is, transitive on its 0-faces) and each of its facets is a regular $(n-1)$-maniplex. Note that, since all polygons (2-polytopes) are regular, for polyhedra (3-polytopes) semiregularity is the same as vertex-transitivity.
Given a maniplex $\mathcal{M}$, we define $r_{i}$ as the flag permutation that maps each flag to its $i$-adjacent flag. It is easy to see that $r_{i}$ is an involution and that, if $|i-j|>1$, then $r_{i}r_{j}$ is also an involution. In particular $r_{i}$ and $r_{j}$ commute whenever $i$ and $j$ are not consecutive. If there is an $i\in[0,n-2]$ such that $(r_{i}r_{i+1})^{2}$ fixes some flag, we say that $\mathcal{M}$ is degenerate, otherwise it is non-degenerate.
@@
If $\mathcal{M}$ is a regular maniplex, for a fixed flag $\Phi$ we can define $\rho_{i}$ as the (unique) automorphism that maps $\Phi$ to $\Phi^{i}$. It is known that $\{\rho_{i}:i\in\{0,1,\ldots,n-1\}\}$ is a generating set for the automorphism group of $\mathcal{M}$ (see [10, Theorem 2B8] for the polytopal case). Moreover, the function mapping $\rho_{i}$ to $r_{i}$ can be extended to a group isomorphism between $\Gamma(\mathcal{M})$ and ${\rm Mon}(\mathcal{M})$ (see, for example, [13, Theorem 3.9] or [20, Section 7]). In particular, the monodromy group of a regular maniplex acts regularly on its flags. It is also worth noticing that a regular maniplex is degenerate if and only if $\rho_{i}\rho_{i+1}$ is an involution for some $i\in[0,n-2]$.
@@ 3 Symmetry type graphs
The symmetry type graph of a polytope (or maniplex) $\mathcal{P}$, denoted by $\mathcal{T}(\mathcal{P})$, is a graph whose vertices are the flag-orbits of $\mathcal{P}$ that has one edge of color $i$ between the orbit of $\Phi$ and the orbit of $\Psi$ if and only if $\Psi$ is in the same orbit as $\Phi^{i}$. If $\Phi$ and $\Phi^{i}$ are in the same orbit, we draw a semi-edge of color $i$ at the vertex representing that orbit in the symmetry type graph.
The symmetry type graph of an $n$-polytope or $n$-maniplex is an example of an $n$-premaniplex, defined as a graph with a coloring of its edges with colors in $\{0,1,\ldots,n-1\}$ such that each vertex is incident to exactly one edge of each color, and satisfying that whenever $|i-j|>1$, the alternating paths of length 4 with colors $i$ and $j$ are closed. For the purposes of this paper all premaniplexes are considered to be connected. In contrast to a maniplex, a premaniplex needs not to be simple. The notions of isomorphism, automorphism, regularity, monodromy, and so on, extend naturallly from maniplexes to premaniplexes.
@@
  • The symmetry type graph of regular $n$-polytopes is denoted by ${\mathbf{1}}^{n}$. It has exactly one vertex and semi-edges of each color from 0 to $n-1$ at that vertex.
@@
  • The symmetry type graph of chiral $n$-polytopes is denoted by ${\mathbf{2}}^{n}_{\emptyset}$: It has exactly two vertices and they are joined by one edge of each color from 0 to $n-1$.
@@
  • More in general, if $\mathcal{P}$ is a polytope with two flag orbits and $I\subset\{0,1,\ldots,n-1\}$ is the set of colors such that $\Phi$ and $\Phi^{i}$ are in the same orbit, then the symmetry type graph of $\mathcal{P}$ is ${\mathbf{2}}^{n}_{I}$, which has exactly two vertices joined by one edge of each color in $\overline{I}$ and with semi-edges of each color in $I$ at each vertex.
@@
The symmetry type graph of a polytope can be used to know how many orbits the polytope has in $i$-faces. In fact, if $\mathcal{X}=\mathcal{T}(\mathcal{P})$ then each connected component of $\mathcal{X}_{\overline{i}}$ (the graph obtained from $\mathcal{X}$ by removing the edges of color $i$), corresponds to an orbit of $i$-faces (see [2, Proposition 3.2]). In particular, $\mathcal{P}$ is $i$-face-transitive if and only if $\mathcal{X}_{\overline{i}}$ (the subgraph of $\mathcal{X}$ induced by edges of color different to $i$) is connected.
@@
A polytope is hereditary if every symmetry of each facet induces a symmetry of the whole polytope, or in other words, if the stabilizer of each facet coincedes with its automorphism group. In terms of the symmetry type graph, a polytope $\mathcal{P}$ with symmetry type graph $\mathcal{X}$ is hereditary if each connected component of $\mathcal{X}_{\overline{n-1}}$ is isomorphic to the symmetry type graph of a facet in the corresponding orbit. In particular, if $\mathcal{P}$ has regular facets and is hereditarty, then each connected component of $\mathcal{X}_{\overline{n-1}}$ must have only one vertex, which in turn means that $\mathcal{X}$ is either ${\mathbf{1}}^{n}$ (implying that $\mathcal{P}$ is regular), or ${\mathbf{2}}^{n}_{\overline{n-1}}$: the graph with two vertices joined by an edge of color $n-1$ and semi-edges of all other colors at each vertex (for more on this, see [11]). Note that if we remove edges of any color other than $n-1$ from these graphs we do not disconnect them, meaning that hereditary polytopes are $i$-face transitive for all $i<n-1$. Polytopes with symmetry type graph ${\mathbf{2}}^{n}_{\overline{n-1}}$ are a particular case of alternating semiregular polytopes: semiregular polytopes with two facet orbits where $(n-1)$-adjacent flags are in different orbits. This kind of polytopes have been studied in [14], for example.
@@
In this paper we want to focus on semiregular polytopes that are the opposite of hereditary: Instead of wanting every symmetry of a facet to induce a symmetry of the whole polytope, we want that none of them do, that is, we want the stabilizer of every facet to be trivial. This means that any two flags in the same facet must be in different orbits. In terms of the symmetry type graph, this means that if $\mathcal{P}$ is a polytope with the desired property and $\mathcal{X}$ is its symmetry type graph, each connected component of $\mathcal{X}_{\overline{n-1}}$ must be isomorphic to a facet in the corresponding orbit.
@@
Note that if a hereditary polytope with regular facets is not regular, it must have exactly two facet orbits. This is not the case when we remove hereditarity: semiregular polytopes can, in principle, have any number of facet orbits.
It is also possible to have a facet-transitive semiregular polytope that is not only not regular, but that has as many flag-orbits as the number of flags in a facet. We will focus most of this paper on constructing examples with this exact characteristic.

### v2

@@ Abstract
Abstract polytopes generalize the face lattice of convex polytopes. A polytope is semiregular if its facets are regular and its automorphism group acts transitively on its vertices. In this paper we construct semiregular, facet-transitive polyhedra with trivial facet stabilizer, showing that semiregular abstract polyhedra can have an unbounded number of flag orbits, while having as little as one facet orbit. We interpret this construction in terms of operations applied to high rank regular and chiral polytopes, and we see how these same operations help us construct alternating semiregular polyhedra (that is, with two facet orbits and adjacent facets in different orbits). Finally, we give an idea to generalize this construction giving examples in higher ranks.
@@ 1 Introduction
Abstract polytopes are combinatorial objects that generalize the face lattice of convex polytopes and other geometric objects. Just as with geometric polytopes, one of the main approaches to study abstract polytopes has been through their symmetry. A flag is typically defined as a set having one face of each rank, and satisfying that all those faces are incident to each other. One can prove that the automorphism group of an abstract polytope acts semiregularly on its flags, and therefore, one can measure how symmetric a polytope is by counting its flag orbits: the fewer flag orbits the more symmetrical. Regular polytopes are those with exactly one flag orbit, and are by far the most studied class of abstract polytope, having a whole book [13] dedicated to them.
@@
When studying other classes of polytopes, some have studied polytopes with a small number of flag orbits (see for example [9], [15], [21], [23]), but some classes where the number of flag orbits is not restricted have been studied too. Such is the case of semiregular polytopes: those with regular facets and where the automorphism group acts transitively on the vertices. Despite the number of flag orbits being unbounded (see [22]), most of the research on semiregular polytopes has been focused on constructions where the result has only one or two flag orbits (see [17], for example). This is due to the fact that the examples constructed are hereditary: every symmetry of a facet induces a symmetry of the whole polytope.
@@
In this paper we construct semiregular polyhedra that are face-transitive and that are the opposite of hereditary: no non-trivial symmetry of a facet induces a symmetry of the whole polyhedron. In other words, our examples have trivial facet stabilizer. This implies that despite having one orbit on vertices and one orbit on faces, the number of orbits on flags in unbounded in our construction.
@@
For our construction it is convenient to think of polytopes as a special class of maniplexes: graphs with a proper coloring of their edges satisfying certain conditions. Thinking of polytopes in terms of graphs allows us to use voltage graphs to construct them, and thanks to the results of [3] and [10], we know conditions on the voltage assignments that tell us when the graphs we obtain represent a polytope.
@@
We will also see that our examples can be interpreted as the result of applying some particular operations to regular and chiral polytopes (those with the highest possible rotational symmetry but no reflections) of high rank. These same operations can be used to obtain alternating semiregular polyhedra with trivial facet stabilizer (that is, semiregular polyhedra with two face-orbits and where each edge lies between faces in different orbits).
@@
In Section˜2 we introduce the basic notions necessary for this paper. Section˜3 presents the concept of symmetry type graph, which is used later in the paper. Section˜4 presents the non-standard definition of graph used in this paper, together with the concept of voltage graphs and how to construct the derived graph of a voltage graph. In Section˜5 we characterize the regular symmetry type graphs that semiregular facet-transitive polyhedra with trivial facet stabilizer may have. Section˜6 and Section˜7 deal with how to construct polyhedra with the symmetry types determined in Section˜5. In Section˜8 we interpret our construction in terms of operations. In Section˜9 we see how to use these operations to get examples of alternating semiregular polyhedra. Finally, in Section˜10 we see how we could generalize some of our constructions to get semiregular, facet-transitive maniplexes with trivial facet stabilizer in higher rank, although we are not sure yet whether they are polytopes or not.
@@ 2 Basic concepts
Abstract polytopes, introduced by Schulte and Danzer in [5], are typically defined as posets with certain properties that make them similar to the face lattice of a convex polytope. However, thanks to [7, Theorem 5.3], we know that they can be equivalently defined as a particular kind of maniplex. For simplicity, we will use this other definition.
An $n$-maniplex (or maniplex of rank $n$) is a connected simple graph $\mathcal{M}$ with a coloring of its edges with colors in $0,n-1]:=\{0,1,\ldots,n-1\}$ such that
@@
  • given two colors $i$ and $j$ such that |ij|>1|i-j|>1, the alternating paths of length 4 using edges of these two colors are closed.
The vertices of a maniplex are referred to as flags. For the purposes of this paper all maniplexes are finite, that is, they have a finite number of flags. Maniplexes are a natural generalization of maps to higher dimensions. To see other definitions of maniplex and also learn the topological interpretation of this definition see [25].
Given an $n$-maniplex $\mathcal{M}$ and a color $i\in[0,n-1]$, we denote by $\mathcal{M}_{\overline{i}}$ the subgraph of $\mathcal{M}$ obtained by removing the edges of color $i$. The connected components of $\mathcal{M}_{\overline{i}}$ are the $i$-faces of $\mathcal{M}$. The $i$-face containing the flag $\Phi$ is denoted by $(\Phi)_{i}$, and the set of all $i$-faces of $\mathcal{M}$ is denoted by $(\mathcal{M})_{i}$. The (n1)(n-1)-faces of an $n$-maniplex are its facets, while the 0-faces and 1-faces are its vertices and edges respectively.
We say that an $i$-face $F$ and a $j$-face $G$ of a maniplex are incident if they have non-empty intersection. If in addition $i\leq j$, we write $F\leq G$. In this way the faces of a maniplex form a poset (see [7, Proposition 3.1]).
In Figure˜1 we see how a hexagon can be interpreted as a 2-maniplex. Note that the vertices of the hexagon correspond to the edges of color 1 of the graph, and the edges of the hexagon correspond to the edges of color 0. Analogously, in Figure˜2 we see a triangular prism interpreted as a 3-maniplex. Here the vertices of the prism correspond to cycles with edges of colors 1 and 2 in the graph, the edges of the prism correspond to cycles with edges of colors 0 and 2 in the graph, and the 2-faces of the prism correspond to cycles with edges of colors 0 and 1 in the graph.
@@
Given a maniplex $\mathcal{M}$ we define its dual $\mathcal{M}^{*}$ as the maniplex we get by replacing the edges of color $i$ with edges of color n1in-1-i for each $i\in[0,n-1]$. Note that the vertices of $\mathcal{M}^{*}$ correspond to the facets of $\mathcal{M}$ and vice-versa. If $\mathcal{M}^{*}$ is isomorphic to $\mathcal{M}$ we say that $\mathcal{M}$ is self-dual.
@@
Path intersection property: Let $\Phi$ and $\Psi$ be two flags of $\mathcal{P}$. If there is a path from $\Phi$ to $\Psi$ that uses only colors in $[0,m]$ and another path from $\Phi$ to $\Psi$ that uses only colors in [k,n1][k,n-1], then there is a path from $\Phi$ to $\Psi$ that uses only colors in $[k,m]$.
In [7, Theorem 5.3] it is proved that this definition is equivalent to the poset definition given by Shculte and Danzer in [5]. From here on, whenever we talk about polytopes in this paper we will be thinking about abstract polytopes.
@@
An isomorphism of $n$-maniplexes is just a graph isomorphism that preserves the colors of the edges. Naturally, an automorphism of a maniplex is just an isomorphism onto itself. The automorphism group of a maniplex $\mathcal{M}$ is denoted by $\Gamma(\mathcal{M})$. It is known that the automorphism group of a maniplex acts freely on its flags (the proof is identical to that of [13, Proposition 2A4], which is the particular case of abstract polytopes). We say that a maniplex $\mathcal{M}$ is regular (some authors prefer reflexible) if $\Gamma(\mathcal{M})$ acts transitively (and therefore regularly) on the flags of $\mathcal{M}$.
A maniplex is called semiregular if it is vertex-transitive (that is, transitive on its 0-faces) and each of its facets is a regular (n1)(n-1)-maniplex. Note that, since all polygons (2-polytopes) are regular, for polyhedra (3-polytopes) semiregularity is the same as vertex-transitivity.
Given a maniplex $\mathcal{M}$, we define $r_{i}$ as the flag permutation that maps each flag to its $i$-adjacent flag. It is easy to see that $r_{i}$ is an involution and that, if |ij|>1|i-j|>1, then $r_{i}r_{j}$ is also an involution. In particular $r_{i}$ and $r_{j}$ commute whenever $i$ and $j$ are not consecutive. If there is an $i\in[0,n-2]$ such that $(r_{i}r_{i+1})^{2}$ fixes some flag, we say that $\mathcal{M}$ is degenerate, otherwise it is non-degenerate.
@@
If $\mathcal{M}$ is a regular maniplex, for a fixed flag $\Phi$ we can define $\rho_{i}$ as the (unique) automorphism that maps $\Phi$ to $\Phi^{i}$. It is known that $\{\rho_{i}:i\in\{0,1,\ldots,n-1\}\}$ is a generating set for the automorphism group of $\mathcal{M}$ (see [13, Theorem 2B8] for the polytopal case). Moreover, the function mapping $\rho_{i}$ to $r_{i}$ can be extended to a group isomorphism between $\Gamma(\mathcal{M})$ and ${\rm Mon}(\mathcal{M})$ (see, for example, [16, Theorem 3.9] or [25, Section 7]). In particular, the monodromy group of a regular maniplex acts regularly on its flags. It is also worth noticing that a regular maniplex is degenerate if and only if $\rho_{i}\rho_{i+1}$ is an involution for some $i\in[0,n-2]$.
@@ 3 Symmetry type graphs
For this paper we need to think of graphs in a way that is not completely standard. A graph $\mathcal{X}$ consists of a set of vertices $V(\mathcal{X})$, a set of darts $D(\mathcal{X})$, a function $I:D(\mathcal{X})\to V(\mathcal{X})$ and an involution $(\cdot)^{-1}:D(\mathcal{X})\to D(\mathcal{X})$. We call $I(d)$ the initial vertex of $d$ and $d^{-1}$ the inverse of $d$. An edge of $\mathcal{X}$ is just a set of the form $\{d,d^{-1}\}$. A semi-edge is a dart $d$ satisfying $d=d^{-1}$. If $d$ is a semi-edge, we abuse language and call the edge $\{d\}$ a semi-edge as well. We think of the edge $e=\{d,d^{-1}\}$ as a line joining the vertices $I(d)$ and $T(d):=I(d^{-1})$ (called the final vertex of $d$), and we think of its darts as the two (or one) possible orientations. A path $W$ on $\mathcal{X}$ is a finite sequence $d_{1}d_{2}\cdots d_{\ell}$ of darts in $\mathcal{X}$ such that $T(d_{i})=I(d_{i+1})$ for all $i\in\{0,1,\ldots,\ell-1\}$. The vertex $I(d_{1})$ is the initial vertex of $W$ while $T(d_{\ell})$ is its final vertex. The number $\ell$ is called the length of $W$. We refer the reader to [12] for details.
The symmetry type graph of a polytope (or maniplex) $\mathcal{P}$, denoted by $\mathcal{T}(\mathcal{P})$, is a graph whose vertices are the flag-orbits of $\mathcal{P}$ and satisfying that it has an edge of color $i$ between the orbit of $\Phi$ and the orbit of $\Psi$ if and only if $\Psi$ is in the same orbit as $\Phi^{i}$. If $\Phi$ and $\Phi^{i}$ are in the same orbit, we draw a semi-edge of color $i$ at the vertex representing that orbit in the symmetry type graph.
The symmetry type graph of an $n$-polytope or $n$-maniplex is an example of an $n$-premaniplex, defined as a graph with a coloring of its edges with colors in $\{0,1,\ldots,n-1\}$ such that each vertex is incident to exactly one edge of each color, and satisfying that whenever |ij|>1|i-j|>1, the alternating paths of length 4 with colors $i$ and $j$ are closed. For the purposes of this paper all premaniplexes are considered to be connected. In contrast to a maniplex, a premaniplex need not be simple. The notions of isomorphism, automorphism, regularity, monodromy, and so on, extend naturally from maniplexes to premaniplexes.
@@
  • The symmetry type graph of regular $n$-polytopes is denoted by ${\mathbf{1}}^{n}$. It has exactly one vertex and semi-edges of each color from 0 to n1n-1 at that vertex.
@@
  • The symmetry type graph of chiral $n$-polytopes is denoted by ${\mathbf{2}}^{n}_{\emptyset}$: It has exactly two vertices and they are joined by one edge of each color from 0 to n1n-1.
@@
  • More generally, if $\mathcal{P}$ is a polytope with two flag orbits and $I\subset\{0,1,\ldots,n-1\}$ is the set of colors such that $\Phi$ and $\Phi^{i}$ are in the same orbit, then the symmetry type graph of $\mathcal{P}$ is ${\mathbf{2}}^{n}_{I}$, which has exactly two vertices joined by one edge of each color in $\overline{I}$ and with semi-edges of each color in $I$ at each vertex (see [4, Section 4]).
@@
The symmetry type graph of a polytope can be used to know how many orbits the polytope has on $i$-faces. In fact, if $\mathcal{X}=\mathcal{T}(\mathcal{P})$ then each connected component of $\mathcal{X}_{\overline{i}}$ (the graph obtained from $\mathcal{X}$ by removing the edges of color $i$), corresponds to an orbit of $i$-faces (see [4, Proposition 3.2]). In particular, $\mathcal{P}$ is $i$-face-transitive if and only if $\mathcal{X}_{\overline{i}}$ is connected.
@@
A polytope is hereditary if every symmetry of each facet induces a symmetry of the whole polytope, or in other words, if the stabilizer of each facet coincides with its automorphism group. In terms of the symmetry type graph, a polytope $\mathcal{P}$ with symmetry type graph $\mathcal{X}$ is hereditary if each connected component of $\mathcal{X}_{\overline{n-1}}$ is isomorphic to the symmetry type graph of a facet in the corresponding orbit. In particular, if $\mathcal{P}$ has regular facets and is hereditary, then each connected component of $\mathcal{X}_{\overline{n-1}}$ must have only one vertex, which in turn means that $\mathcal{X}$ is either ${\mathbf{1}}^{n}$ (implying that $\mathcal{P}$ is regular), or ${\mathbf{2}}^{n}_{\overline{n-1}}$: the graph with two vertices joined by an edge of color n1n-1 and semi-edges of all other colors at each vertex (for more on this, see [14]). Note that if we remove edges of any color other than n1n-1 from these two graphs we do not disconnect them, meaning that hereditary polytopes with regular facets are $i$-face transitive for all i<n1i<n-1. Polytopes with symmetry type graph ${\mathbf{2}}^{n}_{\overline{n-1}}$ are a particular case of alternating semiregular polytopes: semiregular polytopes with two facet orbits where (n1)(n-1)-adjacent flags are in different orbits. This kind of polytope has been studied in [17], for example.
@@
In this paper we want to focus on semiregular polytopes that are the opposite of hereditary: Instead of wanting every symmetry of a facet to induce a symmetry of the whole polytope, we want that none of them do; that is, we want the stabilizer of every facet to be trivial. This means that any two flags in the same facet must be in different orbits. In terms of the symmetry type graph, this means that if $\mathcal{P}$ is a polytope with the desired property and $\mathcal{X}$ is its symmetry type graph, each connected component of $\mathcal{X}_{\overline{n-1}}$ must be isomorphic to a facet in the corresponding orbit.
@@
Note that if a hereditary polytope with regular facets is not regular, it must have exactly two facet orbits. This is not the case when we remove heredity: semiregular polytopes can have any number of facet orbits, even if all facets are isomorphic to each other (see [22, Theorem 5.3]).
In [22], Pisanski, Schulte and Ivić Weiss construct examples of semi-regular $n$-polytopes with trivial vertex stabilizer and show that they can have arbitrarily many flag orbits. However, their examples have (usually) many facet-orbits (even though they are isomorphic to each other). We will show in this article that it is also possible to construct facet-transitive semiregular polytopes that are not only not regular, but that have as many flag-orbits as the number of flags in a facet.